# Asymptotic Inference in a Stationary Quantum Time Series

### v1

@@ 1.1 Introduction
Local asymptotic normality (LAN, Le Cam [LC86]) is a fundamental property of a sequence of statistical experiments, which essentially reduces inference for large sample size to the case of a normal location model. Let $\left(P_{n,\theta},\theta\in\Theta\right)$ be a sequence of families of p.m.’s on $\left(\Omega_{n},\mathcal{A}_{n}\right)$ where $\Theta\subset\mathbb{R}^{k}$; assume that for given $n$, all $P_{n,\theta}$ are mutually absolutely continuous. The sequence is LAN at $\theta\in\mathrm{int}\left(\Theta\right)$ if there exists a positive $k\times k$ matrix $J_{\theta}$ and random $k$-vectors $\Delta_{n,\theta}$ on $\Omega_{n}$ such that $\mathcal{L}\left(\Delta_{n}|P_{n,\theta}\right)\Rightarrow N\left(0,J_{\theta}\right),$ and for $h\in\mathbb{R}^{k}$ one has
@@
When parameters are infinite dimensional, defining a framework of nonparametric inference, the proper analog of LAN to describe risk benchmarks for procedures is asymptotic equivalence in the sense of Le Cam’s $\Delta$-distance. To define it, assume all measurable sample spaces are Polish (complete separable) metric spaces equipped with their Borel sigma algebra. For measures $P,$ $Q$ on the same sample space, let $\left\|P-Q\right\|_{1}$ be $L_{1}$-distance. For the general case where $P,$ $Q$ are not necessarily on the same sample space, suppose $K$ is a Markov kernel such that $KP$ is a measure on the same sample space as $Q$. In that case, $\left\|Q-KP\right\|_{1}$ is defined and will be used to measure the distance between $Q$ and a Markov kernel randomization of $P$.
@@
$\gamma_{j}=\int_{-\pi}^{\pi}\exp\left(-ij\omega\right)f\left(\omega\right)d\omega.$ (1.4)
Define a nonparametric set $\Sigma_{\alpha,M}=B^{\alpha}(M)\cap\mathcal{F}_{M},$ where $B^{\alpha}(M)$ is a (Besov) smoothness class of spectral densities with smoothness coefficient $\alpha$ and $\mathcal{F}_{M}$ is the set of real even positive functions $f$ on $\pi,\pi]$ such that $\left|\log f\right|\leq M$. Then it is shown that observations $y^{(n)}$ with spectral density $f$ are asymptotically equivalent to a white noise model
@@
if the parameter space is given by $f\in$ $\Sigma_{\alpha,M}$ for some $M>0$ and $\alpha>1/2$. This represents the nonparametric (asymptotic equivalence) version of the classical LAN property for parametric models $\left(f_{\theta},\theta\in\Theta\right)$ of spectral densities [Dav73] [Dzh86] [TK00]. Here the Gaussian white noise model (1.5) represents an analog of the basic Gaussian location model (1.2), with the approximation valid globally (over all spectral densities $f\in$ $\Sigma_{\alpha,M}$). Also established were local approximations (via the connection to [GN98]) around a fixed spectral density $f_{0}$ like
@@
which are more suitable for obtaining risk bounds for estimation on $f$ itself, rather than for $\log f$. Here the log-transformation plays the role of a variance stabilizing transformation, removing the factor $f_{0}$ from the noise term and allowing to proceed from the local asymptotic equivalence (1.6) (valid for $f\approx f_{0}$) to the global variant (1.5) (cf. [GN98] for details).
The analog of the $\Delta$-distance for quantum statistical models has been introduced and studied by several authors. In [GK06], [KG09] it was used to define a (strong) quantum analog of the LAN property (1.1) for tensor product models of qubits and finite dimensional states. Alternative approaches to quantum LAN were pursued by [GJ07] and [YFG13], via two different definitions of a quantum likelihood ratio. Recently in [BGN18] the quantum Le Cam distance was used to establish asymptotic equivalence of a tensor product model of infinite dimensional pure states to a quantum Gaussian white noise model. Although the approximation is local, valid in a neighborhood of a fixed pure state (and thus is an analog of (1.6)), it allows to establish some explicit results for nonparametric inference on pure states (estimation and testing).
The object of the present paper is to investigate, with regard to asymptotic equivalence, a quantum Gaussian model studied earlier by Mosonyi [Mos09]. We will consider an $n$-mode quantum Gaussian system to define a quantum Gaussian time series of ”length” $n$.
@@
i.e. the Weyl canonical commutations relations, or CCR. The $C^{\ast}$-algebra generated by $\left\{W\left(x\right),x\in\mathbb{R}^{2n}\right\}$ is the Schrödinger representation of $CCR\left(\mathbb{R}^{2n},D\right)$ ([BR97], 5.2.16). The von Neumann algebra generated by $\left\{W\left(x\right),x\in\mathbb{R}^{2n}\right\}$ is the full algebra $\mathcal{L}(L_{2}\left(\mathbb{R}^{n}\right))$ of bounded operators on $L_{2}\left(\mathbb{R}^{n}\right)$.
@@ 1.3 Shift invariant states
have the same distribution. It is easily seen that this implies shift invariance for the one mode subsystem $\left(Q_{1,}P_{1}\right)$, and also shift invariance for any $r$-mode subsystem $\left(Q_{1,},\ldots,Q_{r,}P_{1},\ldots,P_{r}\right)$, $1\leq r<n$. It follows that the covariance matrix $\Sigma$ is such that all four $r\times r$ submatrices in
@@
are Toeplitz. Equivalently, if $\check{\Sigma}$ is the permutation of $\Sigma\$such that for $\mathbf{\check{R}}:=(Q_{1},P_{1},\ldots,Q_{n},P_{n},)$ we have
@@
then $\check{\Sigma}$ is block Toeplitz, i.e. it is of form $\Sigma=\left(\Sigma_{j-k}^{0}\right)_{j,k=1}^{n}$ where $\left\{\Sigma_{k}^{0}\right\}_{k=0}^{n-1}$ is a sequence of $2\times 2$ matrices. The block Toeplitz structure is familiar in the statistical theory for classical multivariate time series.
@@ 1.4 Gauge invariant states
A state $\rho$ is gauge invariant if for every $z\in\mathbb{C}$, $\left|z\right|=1$ one has $\rho\left(V\left(zu\right)\right)=\rho\left(V\left(u\right)\right),$ $u\in\mathbb{C}^{n}$. A quasifree state $\rho\left(0,\Sigma\right)$ is gauge invariant if and only if
@@
Upon setting $v=iu$, this implies $A\geq I$, and conversely every Hermitian $A\geq I$ satisfies (1.17) and thus is the symbol of an $n$-mode gauge invariant centered Gaussian state. For the gauge invariant centered Gaussian state $\rho=\varphi\left(0,\Sigma\right)$ with symbol $A$, covariance matrix $\Sigma\left(A\right)$ and characteristic function
@@
$\rho=\varphi\left(0,\Sigma\left(A\right)\right)=\mathfrak{N}_{n}\left(0,A\right).$ (1.19)
@@
is the vacuum state for $a=1$ and a thermal state for $a>1$. If $A$ is diagonal $A=\mathrm{diag}\left(a_{1},\ldots,a_{n}\right)>0$ then $\mathfrak{N}_{n}\left(0,A\right)$ is the $n$-fold tensor product of thermal states $\mathfrak{N}_{1}\left(0,a_{j}\right)$.
@@ 1.5 The asymptotic setup
The quantum statistical model for fixed $n$ is now given by a family of gauge invariant and shift invariant centered Gaussian states $\left\{\mathfrak{N}_{n}\left(0,A\right),A\in\mathfrak{A}_{n}\right\}$ where $\mathfrak{A}_{n}$ is a set of $n\times n$ complex Hermitian Toeplitz matrices with $A\geq I$. In accordance wí th the usage in classical statistics, the model might be described as a stationary quantum Gaussian time series. For asymptotic inference in that model, we assume that the $n\times n$ symbols $A=\left(a_{j,k}\right)_{j,k=1}^{n}$ are related to a given positive bounded measurable function $a:\left[-\pi,\pi\right]\rightarrow\mathbb{R}$ as follows:
@@
Here $a_{j}$ are analogs of the autocovariances of a classical stationary complex valued time series, fulfilling $\overline{a_{k}}=a_{-k}$. Accordingly the function $a\left(\omega\right)$ may be described as the quantum spectral density. We assume $a$ to be real and fulfilling $a\geq 1$, and we write $A_{n}\left(a\right)$ for the Hermitian Toeplitz matrix generated by (1.21) for given $a$. We then have $A_{n}\left(a\right)\geq I$ (cf. Lemma 2.10 below); our quantum statistical model is now a family of states
$\left\{\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right),a\in\Theta\right\}$ (1.23)
where $\Theta$ is a family of quantum spectral densities on $\left[-\pi,\pi\right]$ fulfilling $a\geq 1$. Note that if $f$ is a real function with $f\geq 0$ on $\left[-\pi,\pi\right]$ which is even (i.e. symmetric, $f\left(\omega\right)=f\left(-\omega\right)$) then the matrix $A_{n}\left(2\pi f\right)$ is real symmetric nonnegative definite, i.e. it is the covariance matrix of a real random vector. As $A_{n}\left(2\pi f\right)$ is also a sequence of Toeplitz matrices, it would describe the standard setup for asymptotic inference in classical stationary real valued time series [BD91], [Dzh86], [GNZ10]. Indeed comparing (1.21) with (1.4), we see that $a_{j}=\gamma_{j}$ if in (1.21) we set $a\left(\omega\right)=2\pi f\left(\omega\right),$ $\omega\in\left[-\pi,\pi\right]$. Our asymptotic model (1.23), where the spectral density $a$ is the parameter, is a quantum analog of a classical time series, involving the symbol matrices as analogs of covariance matrices. The model has first been treated in [Mos09] in the problem of discrimination between two spectral densities $a_{1},a_{2}$. There the quantum Chernoff bound has been computed for the specified quantum Gaussian models, based on the general form of the quantum Chernoff bound as previously found in [NS09] and [ANSV08].
@@ 1.6 Quantum Le Cam distance
We follow [GJ07] for defining the quantum analog of the $\Delta$-distance (1.3). So far the quantum Gaussian states $\mathfrak{N}_{n}\left(0,A\right)$ have been defined on the von Neumann algebra $\mathcal{L}\left(L_{2}\left(\mathbb{R}^{n}\right)\right)$, but in order to incorporate classical families of probability distributions into this framework, one needs to consider commutative von Neumann algebras defined by spaces $L^{\infty}\left(\mu\right)$ of functions on a $\sigma$-finite measure space $\left(X,\Omega,\mu\right)$. In our appendix section A.1 we clarify how states on a von Neumann algebra $\mathcal{A}$ can be understood as elements of the predual $\mathcal{A}_{\ast}$ of $\mathcal{A}$. The predual $\mathcal{A}_{\ast}$ is a Banach space with norm $\left\|\cdot\right\|_{1}$ such that $\mathcal{A}$ is its dual Banach space, and the states $\varphi$ are positive elements of $\mathcal{A}_{\ast}$ which fulfill $\left\|\varphi\right\|_{1}=1$. In the case $\mathcal{A}=\mathcal{L}\left(L_{2}\left(\mathbb{R}^{n}\right)\right)$, it is well known that a state $\varphi$ has a density operator $\rho_{\varphi}$ (a positive operator on $L_{2}\left(\mathbb{R}^{n}\right)$ with unit trace) such that
@@
is the usual $L^{1}$-distance.
A quantum statistical experiment $\mathcal{E}=\left\{\mathcal{A},\rho_{\theta},\theta\in\Theta\right\}$ is given by a family of states $\rho_{\theta},\theta\in\Theta$ on a von Neumann algebra $\mathcal{A}$ where $\rho_{\theta}\in\mathcal{A}_{\ast}$. As a regularity condition, it is assumed that experiments are homogeneous and in reduced form (cf. Subsection A.1.11). Let $\mathcal{F}:=\left\{\mathcal{B},\sigma_{\theta},\theta\in\Theta\right\}$ be another quantum statistical experiment, indexed by the same parameter $\theta$. The deficiency $\delta\left(\mathcal{E},\mathcal{F}\right)$ is defined as
$\delta\left(\mathcal{E},\mathcal{F}\right):=\inf_{\alpha}\sup_{\theta}\left\|\sigma_{\theta}\circ\alpha-\rho_{\theta}\right\|_{1}$ (1.24)
@@
Consider now sequences of experiments, where the algebras and states depend on $n$, but the parameter space $\Theta$ remains fixed. A sequence $\mathcal{E}_{n}=\left\{\mathcal{A}_{n},\rho_{n,\theta},\theta\in\Theta\right\}$ is said to be asymptotically more informative than $\mathcal{F}_{n}=\left\{\mathcal{B}_{n},\sigma_{n,\theta},\theta\in\Theta\right\}$ if
@@
We write $\mathcal{E}_{n}\succsim\mathcal{F}_{n}$ in this case. If the reverse also holds, i.e. if
@@
As to the statistical meaning of the relation $\mathcal{E}_{n}\succsim\mathcal{F}_{n}$, assume that $\mathcal{E}_{n}=\left\{\mathcal{A}_{n},\rho_{n,\theta},\theta\in\Theta\right\}$ and $\mathcal{F}_{n}=\left\{\mathcal{B}_{n},\sigma_{n,\theta},\theta\in\Theta\right\}$. Then there is a sequence of TP-CP maps (transitions) between preduals $T_{n}:\mathcal{A}_{n\ast}\rightarrow\mathcal{B}_{n\ast}$ such that
@@
Assume that statistical decisions are to be made in the experiment $\mathcal{F}_{n}$. Let $M_{n}$ be an observation channel (measurement) to be applied in $\mathcal{F}_{n}$, such that $M_{n}:\mathcal{B}_{n\ast}\rightarrow L^{1}\left(\nu\right)$ where $\nu$ is a sigma-finite measure on $\left(X,\Omega\right)$. Then $p_{n,\theta}:=M_{n}\left(\sigma_{n,\theta}\right)$ is a $\nu$-probability density on $\left(X,\Omega\right)$, and combining the transitions $M_{n}$ and $T_{n}$, we obtain a $\nu$-probability density $p_{n,\theta}^{\prime}:=M_{n}\left(T_{n}\left(\rho_{n,\theta}\right)\right)$. Then by the contraction property (A.9) of the state transition $M_{n}$
@@
Let a set of $\Omega$-measurable loss functions $W_{n,\theta}:X\rightarrow\left[0,1\right]$, $\theta\in\Theta$ be given. Then a measurement $M_{n}$ as above can be interpreted as a (randomized) decision rule in experiment $\mathcal{F}_{n},$ where the aim is to make $\int W_{n,\theta}p_{n,\theta}d\nu$ small for every $\theta$ (or small in a worst case sense). Then (1.27) implies
@@
In other words, if the sequence $\mathcal{E}_{n}$ is asymptotically more informative than $\mathcal{F}_{n}$ ($\mathcal{E}_{n}\succsim\mathcal{F}_{n}$) then for every randomized decision rule in $\mathcal{F}_{n}$ there exists one in $\mathcal{E}_{n}$ which is asymptotically as good, uniformly in $\theta\in\Theta$.
In applications, when a parameter $\theta$ is to be estimated by $\hat{\theta}_{n}$ , the loss functions $W_{n,\theta}$ are typically of the form $W_{n,\theta}\left(\hat{\theta}_{n}\right)=\ell\left(\eta_{n}\left\|\hat{\theta}_{n}-\theta\right\|\right)$ where $\ell:\left[0,\infty\right)\rightarrow\left[0,1\right]$ is a monotone function and $\eta_{n}\rightarrow\infty$ is a norming sequence. Then the relation $\mathcal{E}_{n}\succsim\mathcal{F}_{n}$ means that estimators in $\mathcal{F}_{n}$ cannot be asymptotically better than those in $\mathcal{E}_{n}$, i. e. the relation provides lower asymptotic risk bounds. If also the converse $\mathcal{F}_{n}\succsim\mathcal{E}_{n}$ can be shown then risk bounds attainable in $\mathcal{E}_{n}$ can also be attained in $\mathcal{F}_{n}$. Applications to optimal estimation of the quantum spectral density will be discussed in a forthcoming version of this preprint.

### v2

@@ 1.1 Introduction
Local asymptotic normality (LAN, Le Cam [LE 86]) is a fundamental property of a sequence of statistical experiments, which essentially reduces inference for large sample size to the case of a normal location model. Let $\left(P_{n,\theta},\theta\in\Theta\right)$ be a sequence of families of p.m.’s on measurable spaces $\left(\Omega_{n},\mathcal{X}_{n}\right)$ where $\Theta\subset\mathbb{R}^{k}$; assume that for given $n$, all $P_{n,\theta}$ are mutually absolutely continuous. The sequence is LAN at $\theta\in\mathrm{int}\left(\Theta\right)$ if there exists a positive $k\times k$ matrix $J_{\theta}$ and random $k$-vectors $\Delta_{n,\theta}$ on $\Omega_{n}$ such that $\mathcal{L}\left(\Delta_{n}|P_{n,\theta}\right)\Longrightarrow_{d}N\left(0,J_{\theta}\right)$ (convergence in distribution), and for $h\in\mathbb{R}^{k}$ one has
@@
When parameters are infinite dimensional, defining a framework of nonparametric inference, the proper analog of LAN to describe risk benchmarks for procedures is asymptotic equivalence in the sense of Le Cam’s $\Delta$-distance. To define it, assume all measurable sample spaces are Polish (complete separable) metric spaces equipped with their Borel sigma-algebra. For measures $P,$ $Q$ on the same sample space, let $\left\|P-Q\right\|_{1}$ be $L_{1}$-distance. For the general case where $P,$ $Q$ are not necessarily on the same sample space, suppose $K$ is a Markov kernel such that $KP$ is a measure on the same sample space as $Q$. In that case, $\left\|Q-KP\right\|_{1}$ is defined and will be used to measure the distance between $Q$ and a Markov kernel randomization of $P$.
@@
$\gamma_{j}=\frac{1}{2\pi}\int_{-\pi}^{\pi}\exp\left(-ij\omega\right)f\left(\omega\right)d\omega.$ (1.4)
Define a function set $\Sigma_{\alpha,M}=B^{\alpha}(M)\cap\mathcal{F}_{M},$ where $B^{\alpha}(M)$ is a Besov-Sobolev smoothness class of spectral densities with smoothness coefficient $\alpha$ and $\mathcal{F}_{M}$ is the set of real even positive functions $f$ on $\pi,\pi]$ such that $\left|\log f\right|\leq M$. Then it is shown that observations $y^{(n)}$ with spectral density $f$ are asymptotically equivalent to a white noise model
@@
if the parameter space is given by $f\in$ $\Sigma_{\alpha,M}$ for some $M>0$ and $\alpha>1/2$. This represents the nonparametric (asymptotic equivalence) version of the classical LAN property for parametric sets $\left(f_{\theta},\theta\in\Theta\right)$ of spectral densities [DAV73] [DZH86] [TK00]. Here the Gaussian white noise model (1.5) represents an analog of the basic Gaussian location model (1.2), with the approximation valid globally (over all spectral densities $f\in$ $\Sigma_{\alpha,M}$). Also established were local approximations (via the connection to [GN98]) around a fixed spectral density $f_{0}$ like
@@
which are more suitable for obtaining risk bounds for estimation on $f$ itself, rather than $\log f$. Here the log-transformation plays the role of a variance stabilizing transformation, removing the factor $f_{0}$ from the noise term and allowing to proceed from the local asymptotic equivalence (1.6) (valid for $f$ close to $f_{0}$) to the global variant (1.5) (cf. [GN98] for details).
The analog of the $\Delta$-distance for quantum statistical models has been introduced and studied by several authors. In [GK06], [KG09] it was used to define a (strong) quantum analog of the LAN property (1.1) for tensor product models of qubits and finite dimensional states. Alternative approaches to quantum LAN were pursued by [GJ07] and [YFG13], via different definitions of a quantum likelihood ratio. In [BGN18] the quantum Le Cam distance was used to establish asymptotic equivalence of a tensor product model of infinite dimensional pure states to a quantum Gaussian white noise model. Although the approximation is local, valid in a neighborhood of a fixed pure state (and thus is an analog of (1.6)), it allows to establish a number of results for nonparametric inference on pure states (estimation and testing).
The object of the present paper is to investigate, with regard to asymptotic equivalence, a quantum Gaussian model studied earlier in [MOS09]. We will consider an $n$-mode quantum Gaussian system to define a quantum Gaussian time series of ”length” $n$.
@@
i.e. the Weyl canonical commutations relations, or CCR. The $C^{\ast}$-algebra generated by $\left\{W\left(x\right),x\in\mathbb{R}^{2n}\right\}$ defines the Schrödinger representation of $CCR\left(\mathbb{R}^{2n},D\right)$ ([BR97], 5.2.16). The von Neumann algebra generated by $\left\{W\left(x\right),x\in\mathbb{R}^{2n}\right\}$ is the full algebra $\mathcal{L}(L_{2}\left(\mathbb{R}^{n}\right))$ of bounded operators on $L_{2}\left(\mathbb{R}^{n}\right)$.
@@ 1.3 Shift invariant states
have the same distribution. It is easily seen that this implies shift invariance for the one mode subsystem $\left(Q_{1},P_{1}\right)$, and also shift invariance for any $r$-mode subsystem $\left(Q_{1,},\ldots,Q_{r,}P_{1},\ldots,P_{r}\right)$, $1\leq r<n$. It follows that the covariance matrix $\Sigma$ is such that all four $n\times n$ submatrices in
@@
are Toeplitz. Equivalently, if $\check{\Sigma}$ is the permutation of $\Sigma\$such that for $\mathbf{\check{R}}:=(Q_{1},P_{1},\ldots,Q_{n},P_{n})$ we have
@@
then $\check{\Sigma}$ is block Toeplitz, i.e. it is of form $\Sigma=\left(\Sigma_{j-k}^{0}\right)_{j,k=1}^{n}$ where $\left\{\Sigma_{k}^{0}\right\}_{k=1-n}^{n-1}$ is a sequence of $2\times 2$ matrices. The block Toeplitz structure is familiar in the statistical theory for classical multivariate time series [TK00].
@@ 1.4 Gauge invariant states
A state $\rho$ is gauge invariant if for every $z\in\mathbb{C}$, $\left|z\right|=1$ one has $\rho\left(V\left(zu\right)\right)=\rho\left(V\left(u\right)\right),$ $u\in\mathbb{C}^{n}$. A quasifree state $\varphi\left(0,\Sigma\right)$ is gauge invariant if and only if
@@
Upon setting $v=iu$, this implies $A\geq I_{n}$, and conversely every $n\times n$ Hermitian matrix $A\geq I_{n}$ satisfies (1.17) and thus is the symbol of an $n$-mode gauge invariant centered Gaussian state. For the gauge invariant centered Gaussian state $\rho=\varphi\left(0,\Sigma\right)$ with symbol $A$, covariance matrix $\Sigma=\Sigma\left(A\right)$ and characteristic function
@@
$\varphi\left(0,\Sigma\left(A\right)\right)=\mathfrak{N}_{n}\left(0,A\right).$ (1.19)
@@
is the vacuum state for $a=1$ and a thermal state for $a>1$. If $A$ is diagonal $A=\mathrm{diag}\left(a_{1},\ldots,a_{n}\right)>I_{n}$ then $\mathfrak{N}_{n}\left(0,A\right)$ is the $n$-fold tensor product of thermal states $\mathfrak{N}_{1}\left(0,a_{j}\right)$.
@@ 1.5 The asymptotic setup
The quantum statistical model for fixed $n$ is now given by a family of gauge invariant and shift invariant centered Gaussian states $\left(\mathfrak{N}_{n}\left(0,A\right),A\in\mathfrak{A}_{n}\right)$ where $\mathfrak{A}_{n}$ is a set of $n\times n$ complex Hermitian Toeplitz matrices with $A\geq I$. In accordance wí th the usage in classical statistics, the model might be described as a stationary quantum Gaussian time series. For asymptotic inference in that model, we assume that the $n\times n$ symbols $A=\left(a_{j,k}\right)_{j,k=1}^{n}$ are related to a given positive bounded measurable function $a:\left[-\pi,\pi\right]\rightarrow\mathbb{R}$ as follows:
@@
Here $a_{k}$ are analogs of the autocovariances of a classical stationary complex valued time series, fulfilling $\overline{a_{k}}=a_{-k}$. Accordingly the function $a\left(\omega\right)$ may be described as the quantum spectral density. We assume $a$ to be real and fulfilling $a\geq 1$, and we write $A_{n}\left(a\right)$ for the $n\times n$ Hermitian Toeplitz matrix generated by (1.21) for given $a$. We then have $A_{n}\left(a\right)\geq I_{n}$ (compare Lemma 2.10 below); our quantum statistical model is now a family of states
$\left(\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right),a\in\Theta\right)$ (1.23)
where $\Theta$ is a family of quantum spectral densities on $\left[-\pi,\pi\right]$ fulfilling $a\geq 1$. Note that if $f$ is a real function with $f\geq 0$ on $\left[-\pi,\pi\right]$ which is even (i.e. symmetric, $f\left(\omega\right)=f\left(-\omega\right)$) then the matrix $A_{n}\left(f\right)$ is real symmetric nonnegative definite, i.e. it is the covariance matrix of a real random vector. As $A_{n}\left(f\right)$ is also a sequence of Toeplitz matrices, it would describe the standard setup for a sequence of covariance matrices in classical stationary real valued time series [BD91], [DZH86], [GNZ10] (except that the standard setup defines the spectral density $f$ without a factor $1/2\pi$ in (1.4)). Indeed comparing (1.21) with (1.4), we see that $a_{j}=\gamma_{j}$ if in (1.21) we set $a\left(\omega\right)=f\left(\omega\right),$ $\omega\in\left[-\pi,\pi\right]$. Our asymptotic model (1.23), where the spectral density $a$ is the parameter, is thus a quantum analog of a classical time series, involving the symbol matrices as analogs of covariance matrices. To our knowledge, the model has first been treated in [MOS09] in the problem of discrimination between two spectral densities $a_{1},a_{2}$. There the quantum Chernoff bound has been computed for the specified quantum Gaussian models, based on the general form of the quantum Chernoff bound as previously found in [NS09] and [ANS+08].
@@ 1.6 Quantum Le Cam distance
We follow [GJ07] for defining the quantum analog of the $\Delta$-distance (1.3). So far the quantum Gaussian states $\mathfrak{N}_{n}\left(0,A\right)$ have been defined on the von Neumann algebra $\mathcal{L}\left(L_{2}\left(\mathbb{R}^{n}\right)\right)$, but in order to incorporate classical families of probability distributions into this framework, one needs to consider commutative von Neumann algebras defined by spaces $L^{\infty}\left(\mu\right)$ of functions on a $\sigma$-finite measure space $\left(\Omega,\mathcal{X},\mu\right)$. In our appendix section A.1 we clarify how states on a von Neumann algebra $\mathcal{A}$ can be understood as elements of the predual $\mathcal{A}_{\ast}$ of $\mathcal{A}$. The predual $\mathcal{A}_{\ast}$ is a Banach space with norm $\left\|\cdot\right\|_{1}$ such that $\mathcal{A}$ is its dual Banach space, and the states $\varphi$ are positive elements of $\mathcal{A}_{\ast}$ which fulfill $\left\|\varphi\right\|_{1}=1$. In the case $\mathcal{A}=\mathcal{L}\left(L_{2}\left(\mathbb{R}^{n}\right)\right)$, it is well known that a state $\varphi$ has a density operator $\rho_{\varphi}$ (a positive operator on $L_{2}\left(\mathbb{R}^{n}\right)$ with unit trace) such that
@@
is the usual $L^{1}$-distance. In general we will refer to $\left\|\cdot\right\|_{1}$ as the predual norm.
A quantum statistical experiment $\mathcal{E}=\left(\mathcal{A},\rho_{\theta},\theta\in\Theta\right)$ is given by a family of states $\rho_{\theta},\theta\in\Theta$ on a von Neumann algebra $\mathcal{A}$ where $\rho_{\theta}\in\mathcal{A}_{\ast}$. As a regularity condition, it is assumed that experiments are homogeneous and in reduced form (cf. Subsection A.1.11). Let $\mathcal{F}:=\left(\mathcal{B},\sigma_{\theta},\theta\in\Theta\right)$ be another quantum statistical experiment, indexed by the same parameter $\theta$. The deficiency $\delta\left(\mathcal{E},\mathcal{F}\right)$ is defined as
$\delta\left(\mathcal{E},\mathcal{F}\right):=\inf_{\alpha}\sup_{\theta}\left\|\rho_{\theta}\circ\alpha-\sigma_{\theta}\right\|_{1}$ (1.24)
@@
Consider now sequences of experiments, where the algebras and states depend on $n$, but the parameter space $\Theta$ remains fixed. A sequence $\mathcal{E}_{n}=\left(\mathcal{A}_{n},\rho_{n,\theta},\theta\in\Theta\right)$ is said to be asymptotically more informative than $\mathcal{F}_{n}=\left(\mathcal{B}_{n},\sigma_{n,\theta},\theta\in\Theta\right)$ if
@@
We write $\mathcal{F}_{n}\precsim\mathcal{E}_{n}$ in this case. If the reverse also holds, i.e. if
$\Delta\left(\mathcal{E}_{n},\mathcal{F}_{n}\right)\rightarrow 0\text{ as }n\rightarrow\infty$
@@
As to the statistical meaning of the relation $\mathcal{F}_{n}\precsim\mathcal{E}_{n}$, it implies there is a sequence of dual channels (TP-CP maps, state transitions) between preduals $T_{n}:\mathcal{A}_{n\ast}\rightarrow\mathcal{B}_{n\ast}$ such that
@@
Assume that statistical decisions are to be made in the experiment $\mathcal{F}_{n}$. Let $M_{n}$ be a dual observation channel (measurement) to be applied in $\mathcal{F}_{n}$, such that $M_{n}:\mathcal{B}_{n\ast}\rightarrow L^{1}\left(\nu\right)$ where $\nu$ is a sigma-finite measure on $\left(\Omega,\mathcal{X}\right)$. Then $p_{n,\theta}:=M_{n}\left(\sigma_{n,\theta}\right)$ is a $\nu$-probability density on $\left(\Omega,\mathcal{X}\right)$, and combining the transitions $M_{n}$ and $T_{n}$, we obtain a $\nu$-probability density $p_{n,\theta}^{\prime}:=M_{n}\left(T_{n}\left(\rho_{n,\theta}\right)\right)$. Then by the contraction property (A.9) of the dual channel $M_{n}$
@@
Let a set of $\mathcal{X}$-measurable loss functions $W_{n,\theta}:\Omega\rightarrow\left[0,1\right]$, $\theta\in\Theta$ be given. Then a measurement $M_{n}$ as above can be interpreted as a (randomized) decision rule in experiment $\mathcal{F}_{n},$ where the aim is to make $\int W_{n,\theta}p_{n,\theta}d\nu$ small for every $\theta$ (or small in a worst case sense). Then (1.27) implies
@@
In other words, if the sequence $\mathcal{E}_{n}$ is asymptotically more informative than $\mathcal{F}_{n}$ ($\mathcal{F}_{n}\precsim\mathcal{E}_{n}$) then for every randomized decision rule in $\mathcal{F}_{n}$ there exists one in $\mathcal{E}_{n}$ which is asymptotically as good, uniformly in $\theta\in\Theta$. Hence risk bounds attainable in $\mathcal{F}_{n}$ can also be attained in $\mathcal{E}_{n}$. Conversely, decision rules in $\mathcal{F}_{n}$ cannot be asymptotically better than those in $\mathcal{E}_{n}$, i. e. the relation provides lower asymptotic risk bounds.